# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

### v1

@@ 4 Achieving Linear Speedup for Push-Pull
$\displaystyle\sum_{i=1}^{m}\Delta_{y}^{(k+i)}=(B-B_{\infty})\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}+(I-B_{\infty})({\mathbf{g}}^{(k+m)}-{\mathbf{g}}^{(k)}).$ (6)
Equation (6) reveals a contraction-type recursion governing the sum of consecutive error terms. Informally, if we restrict our analysis to terms involving $\sigma^{2}$, we observe that $\|{\mathbf{g}}^{(k+m)}-{\mathbf{g}}^{(k)}\|F^{2}={\mathcal{O}}(\sigma^{2})$. This contraction property further implies that the aggregated term $\|\sum{i=1}^{m}\Delta_{y}^{(k+i)}\|_{F}^{2}$ maintains the same ${\mathcal{O}}(\sigma^{2})$ scaling. Consequently, we can deduce that individual terms satisfy $\|\Delta_{y}^{(k)}\|_{F}^{2}={\mathcal{O}}(\frac{\sigma^{2}}{m})$.
This scaling relationship becomes particularly significant when $m$ is sufficiently large: in this regime, the $\sigma^{2}$-dependent component of the extra error term becomes negligible. As a result, our analysis preserves the crucial linear speedup property. To formalize this logic, we start from the following contraction relationship.
@@
For any integers $k\geq 0$ and $m\geq 1$, we have
$\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\pi_{A}^{\top}\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right){\mathbf{y}}^{(k)}+\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (7)
Furthermore, we can define a ”noiseless” version for the right-hand side, i.e.
$H_{m,k}:=\pi_{A}^{\top}\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right){\mathbf{y}}^{(k)}+\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})),$
and the distance therein can be bounded by:
$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\varepsilon^{(k+i)}-H_{m,k}\|_{F}^{2}]\leq 2ns_{B}^{2}\sigma^{2}.$ (8)
The proof of Lemma 3 is provided in Appendix B.1. This lemma implies that the summation of the $m$ error terms can be represented as $H_{m,k}$ plus a bounded noise term. Moreover, the term $H_{m,k}$ exclusively involves difference of deterministic gradients and a single instance of ${\mathbf{y}}^{(k)}$, making it straightforward to estimate its magnitude.
To effectively leverage the result of Lemma 3, we introduce an $m$-step descent lemma. This new lemma follows by an estimation on $H_{m,k}$ and applying the standard $L$-smoothness descent lemma to the following recursion:
@@
Suppose Assumption 34 and Proposition 1 hold. For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$. The following inequality holds for any integers $k\geq 0,m\geq 1$:
$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c(1-4\alpha mcL)}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (10)
@@
$\displaystyle\quad+\frac{L}{16m}\Delta_{1}^{(k)}+\frac{L}{16m}\Delta_{2}^{(k)}+{3\alpha^{2}Ls_{B}^{2}}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}],$
@@
$\displaystyle\Delta_{1}^{(k)}=\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}=\sum_{j=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
Remarkably, if we divide both sides of (10) by $\frac{c\alpha m}{4}$, the first noise term simplifies to $\frac{8\alpha cL}{n}\sigma^{2}$, which corresponds to the standard linear speedup term. The second noise term simplifies to $\frac{8\alpha ns_{B}^{2}}{cm}\sigma^{2}$. When we select some $m$ larger than $c^{-2}n^{2}s_{B}^{2}$, the second noise term should have a smaller order than the first term, thereby ensuring the achievement of the standard linear speedup, which verifies our logic before. Next, we try to estimate $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$.
Lemma 5 (Consensus Error).
Suppose Assumption 34 and Proposition 1 hold. With $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$ defined above, when $\alpha\leq\min\{\frac{3c}{16mL},\frac{1}{6s_{B}L}\}$, we have
$\displaystyle\quad\frac{L}{16m}\mathbb{E}_{k}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]\leq\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}.$
where $\Delta_{x}:={\mathbf{x}}^{(k)}-A_{\infty}{\mathbf{x}}^{(k)}$, $\Delta_{y}^{(k)}:={\mathbf{y}}^{(k)}-B_{\infty}{\mathbf{y}}^{(k)}$.
Lemma 5 transfers the complicated error terms to commonly-seen consensus error terms. Now, we can apply techniques similar to [38] to bound $\Delta_{x}$ and $\Delta_{y}$.
Lemma 6.
Suppose Assumption 34 and Proposition 1 hold. For any $T\geq 1$, when $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, we have
$\displaystyle\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k+1)}\|^{2}]$ $\displaystyle\leq 2\alpha^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{y}^{(k)}\|^{2}]+2\alpha^{2}n^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|^{2}]+2n\alpha^{2}s_{A}^{2}(T+1)\sigma^{2},$
and
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}].$
Now we are ready to state the main result of this paper.
@@
Suppose Assumption 34 and Proposition 1 hold. For given positive integers $m$ and $K$, we define $[m^{*}]:=\{0,m,\ldots,(K-1)m\}$ and $T:=mK$. Let $m=\lceil 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})\rceil$. When the learning rate $\alpha$ satisfies $\alpha\leq\min\{\frac{1}{8cL},\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L}\}$, we have
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{30c\alpha L}{n}\sigma^{2}+\frac{64n\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}.$
where $C_{1}$ is a constant determined by $A$ and $B$. Furthermore, if we choose
$\alpha=\min\left\{\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L},\frac{1}{8cL},\frac{\sqrt{n\Delta}\sigma}{3c\sqrt{LT}}\right\},$
we will obtain that
$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{C_{\Delta}L\Delta}{T},$ (11)
where
$C_{\Delta}=32+32ncs_{A}+84n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$
is a constant decided by $A$ and $B$.
Comparison with existing analyses. Theorem 1 provides the first finite-time non-convex convergence result for Push-Pull, while previous analyses focus either on the strongly convex case [45, 40] or on variants of Push-Pull relying on specific topologies [5]. The first term $\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ explicitly shows that Push-Pull enjoys a linear speedup on arbitrary strongly connected topologies.
Impact of $m$. In our analysis, we consider a $m$-step gradient descent. The number $m$ effectively controls the $\sigma^{2}$-related term in the extra error. A common question is can we choose an arbitrarily large $m$ in this analysis. We can do that, but choosing a larger $m$ will force us to select a smaller learning rate, which finally leads to a larger transient time. To sum up, number $m$ provides a chance for tradeoff between the noise level and the transient time.
Transient time. Our theorem provides a coarse estimate for the transient time, which is of ${\mathcal{O}}(nC_{\Delta}^{2})$. We do not recommend using this to estimate the real transient time, but it still reflects the impact of topology. For example, if we use a directed ring graph where $s_{A}=s_{B}={\mathcal{O}}(n)$, $c={\mathcal{O}}(1)$, the transient time can be ${\mathcal{O}}(n^{17})$. If we use a directed exponential graph [3] where $s_{A}=s_{B}={\mathcal{O}}(1)$, $c={\mathcal{O}}(1)$, the transient time will be ${\mathcal{O}}(n^{5})$.
@@ Appendix A Rolling-Sum Lemma
$\displaystyle s_{A}:=\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}.$
where $\beta_{A}$, $\kappa_{A}$ are defined in Proposition 2. The row-stochastic matrix $A$ and the subscript $A$ can be replaced by a column-stochastic matrix $B$ without changing the lemma.
@@
$\displaystyle s_{A}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}.$ (14)
@@
$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}(x+p+\frac{\beta_{A}^{x}}{1-\beta_{A}})\leq M_{A}(p+\frac{1}{1-\beta_{A}}).$
By the definition of $p$, we have $p\leq\frac{\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}$. This finishes the proof of (14).
@@ B.1 Proof of Lemma 3
For $1\displaystyle 1\leq i\leq m-$ 1, consider
$\displaystyle(I-B_{\infty}){\mathbf{y}}^{(k+i)}$ $\displaystyle=(I-B_{\infty})(B{\mathbf{y}}^{(k+i-1)}+{\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k+i-1)})$
$\displaystyle=(B-B_{\infty}){\mathbf{y}}^{(k+i-1)}+(I-B_{\infty}){\mathbf{g}}^{(k+i)}-(I-B_{\infty}){\mathbf{g}}^{(k+i-1)}$
Sum it up for $1\displaystyle 1\leq i\leq m-$ 1, we obtain that
$(I-B_{\infty})\sum_{i=1}^{m-1}{\mathbf{y}}^{(k+i)}=(B-B_{\infty})\sum_{i=0}^{m-2}{\mathbf{y}}^{(k+i)}+(I-B_{\infty}){\mathbf{g}}^{(k+m-1)}-(I-B_{\infty}){\mathbf{g}}^{(k)}.$
Similarly, consider $(B-B_{\infty}){\mathbf{y}}^{(k+i)}$, $2\displaystyle 1\leq i\leq m-$ 2, we have
$(B-B_{\infty})\sum_{i=1}^{m-2}{\mathbf{y}}^{(k+i)}=(B^{2}-B_{\infty})\sum_{i=0}^{m-3}{\mathbf{y}}^{(k+i)}+(B-B_{\infty}){\mathbf{g}}^{(k+m-2)}-(B-B_{\infty}){\mathbf{g}}^{(k)}.$
Repeat this procedure for $m-3,m-4,\ldots,1$ and sum them up, we finally arrive at
$\sum_{i=0}^{m-1}(I-B_{\infty}){\mathbf{y}}^{(k+i)}=\sum_{i=0}^{m-1}(B^{i}-B_{\infty}){\mathbf{y}}^{(k)}+\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)}),$
which is just the first statement. Next, note that the difference between $\sum_{i=0}^{m-1}\varepsilon^{(k+i)}$ and $H_{m,k}$ equals to:
$\pi_{A}^{\top}\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))-\pi_{A}^{\top}\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)})).$
By taking the squared norm, we obtain that:
$\displaystyle\quad\sum_{i=1}^{m-1}\|\pi_{A}^{\top}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|^{2}$
$\displaystyle+\|\pi_{A}^{\top}(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|^{2}+\text{Inner Product Terms}.$
When we take the expectation conditioned on ${\mathcal{F}}_{k}$, all the inner product terms become zero. For example, for any $1\displaystyle 0\leq i<j\leq m-$ 1,
$\displaystyle\mathbb{E}_{k}\left[\left\langle(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})),(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\right\rangle\right]$
$\displaystyle=\mathbb{E}_{k}[\mathbb{E}_{k+j}[\text{This inner product}]]$
$\displaystyle=\mathbb{E}_{k}\left[\left\langle(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})),0\right\rangle\right]=\mathbb{E}[0]=0.$
Therefore, we only need to estimate the squared terms. By the Cauchy-Schwarz inequality, and $\|\pi_{A}\|\leq 1$, we have
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\pi_{A}^{\top}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|^{2}]$
$\displaystyle\leq\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\pi_{A}\|^{2}\|(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|_{F}^{2}]$
$\displaystyle\leq\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2}^{2}\mathbb{E}_{k}[\|{\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})\|_{F}^{2}]$
$\displaystyle\leq n\sigma^{2}(\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2})^{2}\leq ns_{B}^{2}\sigma^{2},$
and
$\displaystyle\quad\mathbb{E}_{k}[\|\pi_{A}^{\top}(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|^{2}]$
$\displaystyle\leq\mathbb{E}_{k}[\|\pi_{A}\|^{2}\|(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq\|\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})\|_{2}^{2}\mathbb{E}_{k}[\|{\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq(\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2})^{2}n\sigma^{2}\leq ns_{B}^{2}\sigma^{2}.$
By combining the two parts, we finish the proof. $\square$
B.2 Proof of Lemma 4
@@
$\mathbb{E}[\bar{g}^{(k+i)}]=\mathbb{E}[\mathbb{E}_{k}[\bar{g}^{(k+i)}]]=\mathbb{E}[\overline{\nabla f}^{(k+i)}],\quad\forall i\geq 0.$
@@
By taking expectation (from time 00 0), we obtain that
$\displaystyle\mathbb{E}[f(\hat{x}^{(k+m)})-f(\hat{x}^{(k)})]\leq-c\alpha\sum_{i=0}^{m-1}\mathbb{E}\left[\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (18)
$\displaystyle\quad-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]$
$\displaystyle\quad+\alpha^{2}L\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
@@
$\mathbb{E}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]\leq n^{-1}m\sigma^{2}+\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}],$
where we use the fact that $\bar{g}^{(k+i)}$ are conditional unbiased estimate and
$\mathbb{E}\left[\left\langle\bar{g}^{(k+i)}-\overline{\nabla f}^{(k+i)},\bar{g}^{(k+j)}-\overline{\nabla f}^{(k+j)}\right\rangle\right]=0.$

### v2

@@ 4 Achieving Linear Speedup for Push-Pull
$\displaystyle\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (11)
To understand this equation, we propose the following important supporting lemma:
@@
For any $l\geq 0$ and any matrices $M^{(0)},M^{(1)},\ldots,M^{(l)}\in\mathbb{R}^{n\times n}$. For any $k\geq 0$, If $H=\sum_{i=0}^{l}M^{(i)}{\mathbf{g}}^{(k+i)}$ is a linear combination of stochastic gradients, then
$\displaystyle\mathbb{E}_{k}[\|H\|_{F}^{2}]\leq n\sigma^{2}\sum_{i=0}^{l}\|M^{(i)}\|_{2}^{2}+\mathbb{E}_{k}[\|\sum_{i=0}^{l}M^{(i)}\nabla f({\mathbf{x}}^{(k+i)})\|_{F}^{2}].$
Proof.
This is a natural corollary of Assumption 3. ∎
Now we turn to the noise estimation for $\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$. A common analysis using Cauchy’s inequality typically provides
$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]\leq m\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\Delta_{y}^{(k+i)}\|_{F}^{2}]\lesssim m^{2}\sigma^{2}.$ (12)
Here, we use $\lesssim$ to absorb matrix-related constants and focus on $\sigma^{2}$-related terms. Nevertheless, analysis using Equation (11) and Lemma 3 can be significantly better. Proposition 2 ensures that $\textstyle\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}+\|\sum_{j=1}^{m-1}B^{m-j-1}-B_{\infty}\|_{2}^{2}$ is finite. Therefore, applying Lemma 3 to Equation (11) tells us
$\displaystyle\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]\lesssim\mathbb{E}_{k}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+\sigma^{2}$ (13)
$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
$\displaystyle\lesssim\sigma^{2}+\alpha^{2}L^{2}m^{2}\sigma^{2}.$
Estimate from (13) sufficiently exploits the structure of gradient tracking, thus exhibiting ${\mathcal{O}}(m^{2})$ improvement compared to a naive estimate (12). To effectively leverage the result of  (13), we introduce an $m$-step descent lemma. It comes from applying the standard descent lemma to the following recursion:
@@
Suppose Assumption 34 and Proposition 1 hold. For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$. When $\alpha\leq\frac{1}{6cmL}$, the following inequality holds for any integers $k\geq 0,m\geq 6c^{-2}s_{B}^{2}$:
$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (15)
@@
$\displaystyle\quad+\frac{c\alpha L^{2}}{2}\Delta_{1}^{(k)}+\frac{c\alpha L^{2}}{2}\Delta_{2}^{(k)}+\frac{3\alpha s_{B}^{2}}{cm}\|\Delta_{y}^{(k)}\|_{F}^{2},$
@@
$\displaystyle\Delta_{1}^{(k)}=\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}=\sum_{j=1}^{m}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
The remaining errors are estimated in the next lemma.
Lemma 5 (Remaining Errors).
Suppose Assumption 34 and Proposition 1 hold. With $\Delta_{1}^{(k)}$, $\Delta_{2}^{(k)}$ defined above, when $\alpha\leq\min\{\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\}$ and $m\geq C_{m}$, we have
$\displaystyle\quad\frac{c\alpha L^{2}}{2}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{3\alpha s_{B}^{2}}{cm}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq\frac{\alpha c}{4}\sum_{t=0}^{mK-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(t)}\|^{2}]+\frac{\alpha cm}{4}(\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}})\sigma^{2}+C_{\Delta,1}\frac{c\alpha L\Delta}{4m}.$
where we define $[m^{*}]:=\{0,m,\ldots,(K-1)m\}$, and $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants. $s_{B^{m}}$ is the measure defined for $B^{m}$, see (3). Generally, we have $s_{B^{m}}\leq s_{B}$, and $s_{B^{m}}$ converges to $\|I-B_{\infty}\|_{2}$ for large $m$.
We are ready to state the main result of this paper.
@@
Suppose Assumption 34 and Proposition 1 hold. For given positive integers $m$ and $K$, we define $T:=mK$. When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq C_{m}$, we have
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{100c\alpha L}{n}\sigma^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{mT}.$ (16)
where $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants. Furthermore, if we choose
$\displaystyle m$ $\displaystyle=\lceil\max\{4c^{-1}s_{B}^{1.5}s_{B^{m}}\left(\frac{nT\sigma^{2}}{L\Delta}\right)^{0.25},C_{m}\}\rceil,$
$\displaystyle\alpha$ $\displaystyle=\min\left\{\frac{\sqrt{n\Delta}\sigma}{10c\sqrt{LT}},\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\right\},$
we reach that
$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]$ $\displaystyle\leq\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{L\Delta(m^{-1}C_{\Delta,1}+10m\sqrt{s_{B}s_{B^{m}}^{2}}+c^{-1})}{T}$ (17)
$\displaystyle=\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)^{\frac{3}{4}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right).$
Proof.
See Appendix B.3
Comparison with existing analyses. Theorem 1 provides the first finite-time non-convex convergence result for stochastic Push-Pull, while previous analyses focus either on the strongly convex case [46, 40] or on variants of Push-Pull relying on specific topologies [5]. The first term $\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ explicitly shows that Push-Pull enjoys a linear speedup on arbitrary topologies satisfying Proposition 2.
Impact of $m$: In our analysis, we consider $m$-step gradient descent. The parameter $m$ effectively controls the terms related to $\sigma^{2}$ from additional errors, as illustrated by the $\mathcal{O}(m^{-2}\sigma^{2})$ term in (16). $m$ cannot be arbitrarily large because larger values of $m$ restrict the range of the learning rate, resulting in slower convergence and longer transient times in the analysis. In summary, the parameter $m$ allows for a tradeoff between noise level and transient time.
Transient time. Our theorem is able to provide a coarse estimate for the transient time, which is
$\max\{n^{2}c^{-4}s_{B}^{8},n^{7}c^{-4}s_{A}^{2}s_{B}^{3},nc^{-4}s_{B}^{4}\}.$
The order of $n$ is overestimated because we frequently use $\|\pi_{A}\|\leq 1$ and $\|\pi_{B}\|\leq 1$, leading to a loss of $n$ when $A$ or $B$ is doubly-stochastic.
@@ Appendix A Rolling-Sum Lemma
$\displaystyle s_{A}:=\max\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\}\leq\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$
where $\beta_{A}$, $\kappa_{A},M_{A}$ are defined in Proposition 2. The row-stochastic matrix $A$ and the subscript $A$ can be replaced by a column-stochastic matrix $B$ without changing the lemma.
@@
$\displaystyle s_{A}\leq\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$ (20)
@@
$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}(x+p+\frac{\beta_{A}^{x}}{1-\beta_{A}})\leq M_{A}(p+\frac{1}{1-\beta_{A}})=\frac{M_{A}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$
By the definition of $p$, we have $p\leq\frac{\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}$. This completes the estimation of $\sum_{i=0}^{\infty}\|A^{k-i}-A_{\infty}\|_{2}$. For the summation of the squared norm, we can use the same $p$ and obtain the upper bound without modifying the previous steps:
$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}^{2}\leq M_{A}^{2}(p+\frac{1}{1-\beta_{A}})=\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$
This finishes the proof of (20).
@@ B.1 Proof of Lemma 4
$\mathbb{E}_{k}[\bar{g}^{(k+i)}]=\mathbb{E}_{k}[\mathbb{E}_{k+i}[\bar{g}^{(k+i)}]]=\mathbb{E}_{k}[\overline{\nabla f}^{(k+i)}],\quad\forall i\geq 0.$
@@
By taking expectation conditioned on ${\mathcal{F}}_{k}$, we obtain that
$\displaystyle\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]-f(\hat{x}^{(k)})\leq-c\alpha\sum_{i=0}^{m-1}\mathbb{E}_{k}\left[\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (24)
$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+c^{2}\alpha^{2}L\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]$
$\displaystyle\quad+\alpha^{2}L\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
@@
$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]\leq n^{-1}m\sigma^{2}+\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}],$
where we use the fact that $\bar{g}^{(k+i)}$ are conditionally unbiased estimates and
$\mathbb{E}_{k}\left[\left\langle\bar{g}^{(k+i)}-\overline{\nabla f}^{(k+i)},\bar{g}^{(k+j)}-\overline{\nabla f}^{(k+j)}\right\rangle\right]=0.$